Stock X has a beta of 1.6 and an expected return of 18.2 percent. Stock Y has a beta of 1.19 and an expected return of 15.49 percent. What is the risk-free rate if these securities both plot on the security market line?
Answer #1
Per CAPM, we have
18.20 = rf+1.60*p ------Eq 1 and
15.49 = rf+1.19*p ------Eq 2
where, rf = risk free rate and p = market risk premium
Solving for p,
Subtracting Eq 2 from Eq 1
2.71 = 0.41*p
p = 2.71/0.41 = 6.61
Now from Eq 1
18.20 = rf+1.60*6.61
rf = 18.20-(1.60*6.61) = 7.62
Answer: risk free rate = 7.62%
Check:
7.62+1.19*6.61 = 15.49%